OpenStudy (anonymous):

given that f(x)=2x^2-5. find the values of f(0), f(3), f(10).

Parth (parthkohli):

\[f(0) = 2(0)^2 - 5\]\[f(1) = 2(1)^2 - 5\]\[f(2) = 2(2)^2 - 5 \\ . \\ . \\ .\]

OpenStudy (anonymous):

For f(0).. put x= 0 in the Eq

Parth (parthkohli):

\[f(\color{#C00}{x}) = 2\color{#C00}{x}^2 - 5 \\ f(\color{#C00}{0}) = 2(\color{#C00}{0})^2 - 5 \\ . \\ . \\ .\]